Lexis co. has a net income for $156 on total sales of $1120. cost are $730 and depreciation is $150. the tax rate is 35 percent. the firm does not have interest expenses. how much is the operating cash flow?
Answer #1

Operating Cash Flow = (Sales - Costs)*(1-tax) + Tax*depreciation

= (1120-730)*(1-0.35) + 0.35*150

= 390*0.65 + 0.35*150

= 253.5 + 52.5 = $306